OpenStudy (anonymous):

Medal Bias binding is to be sewed around the edge of a rectangular tablecloth measuring 72 in by 45 in. If the bias binding comes in package containing 15 ft of binding, how many packages of bias binding are needed for the tablecloth?

OpenStudy (anonymous):

\[\frac{2 (72+45)}{12\ 15}=\frac{13}{10} \] 2 packages.

OpenStudy (anonymous):

so is the answer 13/10